Weekly LIPAc highlights (22/Febuary-26/February)

Global Alignment survey

Additional alignment survey has been completed and the results shared and thoroughly discussed. The realignment of the second solenoid magnet (SOL2) of the LEBT (Low Energy Beam Transport) to the entrance of the RFQ will be performed starting next week. The correction of the alignment between the end of the RFQ and the beginning of the MEBT (Medium Energy Beam Transport) will be performed in the next maintenance period and before starting high duty beam operation. Correction of the mismatch between RFQ and MEBT, considering the size of the RFQ, is challenging and laborious (detailed procedures will be prepared).

Radio Frequency (RF) System and RFQ cavity                                                             

The start-up of the RF system with the calibration of the LLRF (Low Level RF) and the RF chains continued. The installation of Low Pass Filters into the PSYS-LLRF (Protection System – Low Level RF) interface boards was completed. During first RF injections @ 132 kV, D+ level, noise issues on PSYS interlock and water flow meter appear when pulse length exceeds 250us and are under investigation. Check-out of the RFQ cavity with RF injection is pending the completions of these activities.

Beam Dump

As part of the BD skid commissioning, the update of the PLC software for the BD cooling skid is on-going with remote support by CIEMAT.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.